Exponential equation without logs. Solve for x:
5^(2x+4)= 125^(x-4)

Answer by stanbon(57290)

Exponential equation without logs. Solve for x:
5^(2x+4)= 125^(x-4)
5^(2x+4) = [5^3]^(x-4)
5^(2x+4) = 5^(3x-12)
Since the bases are the same the exponents are equal.
2x+4 = 3x-12
x = 16
============
Cheers,
Stan H.